# Anomalous current fluctuations from Euler hydrodynamics

### v2

@@ Abstract
The assertion that such fluctuations are always transported by the Euler equations regardless of the dynamical exponent stems from the observation that the only source of typical current fluctuations in the SCCA are the interactions of ballistically-propagating initial typical fluctuations. A similar observation was recently also made for Dirac fluids in Ref. [18]. This is in fact in the same spirit as diffusion from convection, which is known to be the only contribution to the diffusion constants in integrable systems [30, 31, 32] (we show that this is also the case in the SCCA in SM) and is generically also present even in non-integrable systems, provided that the eigenvalues of the flux jacobian $\mathsf{A}$ are non-degenerate [33].
To illustrate the idea, we focus on the case of half-filling $b=0$ and leave the details and $b\neq 0$ for the SM. Since in this case the dynamical exponent of charge transport is $z=2$, the fluctuating fields $\varrho_{i}(x,t)$ need to be rescaled to the diffusive coordinates $\varrho_{i}(x,t)=\hat{\varrho}_{i}(\sqrt{\tau}x,\tau t)$ with the associated currents $j_{i}(x,t)=\sqrt{\tau}\hat{\jmath}_{i}(\sqrt{\tau}x,\tau t)$. The fields $\varrho_{i}(x,t)$ then fluctuate initially as $\varrho_{i}(x,0)\simeq\rho_{i}+\tau^{-1/4}\delta\varrho_{i}(x,0)$, where $\delta\varrho_{i}(x,0)$ is Gaussian distributed as $\langle\delta\varrho_{i}(x,0)\delta\varrho_{j}(y,0)\rangle=\mathsf{C}_{ij}\delta(x-y)$ with the susceptibility matrix $\mathsf{C}_{ij}=\partial\rho_{i}/\partial\beta^{j}$ (see SM for the explicit expression). The fluctuating particle fields $\delta\varrho_{\pm}(x,0)$ are then transported by the Euler equation $\partial_{t}\delta\varrho_{\pm}(x,t)\pm\sqrt{\tau}\partial_{x}\delta\varrho_{\pm}(x,t)=0$, which can be solved as $\delta\varrho_{\pm}(x,t)\simeq\rho_{\pm}+\tau^{-1/4}\delta\varrho_{\pm}(x\mp\sqrt{\tau}t,0)$. Likewise the charge field $\delta\varrho_{c}(x,t)$ also fluctuates as $\delta\varrho_{c}(x,t)\simeq\tau^{-1/4}\delta\varrho_{c}(x-X_{d}(t),0)$ where $X_{d}(t)$ is the characteristics that satisfies ${\rm d}X_{d}(t)/{\rm d}t=\sqrt{\tau}v(X_{d}(t),t)$. With these we obtain the leading behavior of the fluctuating charge current at the origin $j_{c}(0,t)\simeq\tau^{-1/4}\partial_{t}\int_{0}^{X_{d}(t)}{\rm d}x\,\delta\varrho_{c}(-x,0)$, which in turn allows us to write down the functional integral that describes the contribution to the typical fluctuations by these fields
@@
where $\mathsf{C}^{ij}=[\mathsf{C}^{-1}]_{ij}$ and $Z$ is the normalization factor of the initial probability distribution. This integral can be performed exactly (see SM), yielding the non-Gaussian distribution
$\mathcal{P}_{\mathrm{typ}}(j)=\frac{1}{2\pi\Delta}\int_{\mathbb{R}}\frac{{\rm d}u}{\sqrt{|u|}}e^{-\frac{u^{2}}{2\Delta^{2}}-\frac{j^{2}}{2|u|}},$ (19)
which matches the result reported in Ref. [10]. The breakdown of Gaussianity in Eq. (19) can be understood by noting which fluctuating fields are responsible for charge current fluctuations. At half-filling, it follows from Eq. (18) that charge fluctuations $\delta\varrho_{c}(x,t)$ that travel along the characteristics $X_{d}(t)$ and cross $x=0$ between $t=0$ and $t=\tau$ contribute to current fluctuations. Since charge fluctuations are transported non-linearly, Gaussianity of inital fluctuations is no longer preserved and typical fluctuation become non-Gaussian. The situation changes dramatically away from half-filling, i.e. for $b\neq 0$ where only particle fluctuations $\delta\varrho_{\pm}(x,t)$ emanating along light cones contribute to charge current fluctuations. Since particle fluctuations are transported linearly, Gaussianity of initial fluctuations is preserved during the dynamics, giving rise to Gaussian typical fluctuations (see SM for details).
Typical fluctuations of stochastic dynamics.—The distribution (19) features a cusp at the origin. Numerical simulations for $\Gamma>0$ at half-filling indicate that the cusp is rounded off with stochastic crossings while the distribution remains non-Gaussian. This motivates a generalization of the single-file distribution (19) of the form
@@
which reduces to the single-file distribution (19) as $\omega\to 0$ and to a Gaussian distribution (with variance $\omega$) as $\omega\to\infty$. Treating $\omega$ and $\sigma$ as $\Gamma$-dependent fitting parameters, the distribution (20) accurately captures the asymptotic typical distribution for $\Gamma\geq 0$ across 5 order of magnitude as shown in Fig.  2.
@@
Conclusions.—We have studied anomalous charge current fluctuations in a family of stochastic cellular automata and characterized both large and typical fluctuations based only on Euler hydrodynamics, reproducing the microscopic results obtained in Ref. [10] when specializing to the single-file (i.e. deterministic) limit $\Gamma=0$. The key insight is that initial fluctuations, carried by Euler hydrodynamics, fully control fluctuations both on the typical and large scales in the single-file case. To implement the idea, we employed BMFT to compute the SCGF and proposed a BMFT-based approach to describe typical fluctuations induced by initial Gaussian fluctuations (in a forthcoming work Ref. [34], it is also shown how BMFT provides the full diffusive-order hydrodynamic equation, where the long-range correlations predicted by BMFT are the source of diffusive-order terms). Following Ref. [18], our work further clarifies the hydrodynamic origin of anomalous fluctuations, which we believe is generic among systems that exhibit anomalous fluctuations.
@@
While initial fluctuations remain the only source of large fluctuations even in the stochastic case $\Gamma\neq 0$, it turns out that typical fluctuations are nontrivially influenced by the presence of stochasticity, which is not prescribed by initial fluctuations. We expect that stochasticity induces normal diffusion, and it would be very interesting to reproduce Eq. (20) by extending the formalism we introduced to incorporate normal diffusion. The observation of the distribution (19) in a classical spin chain [8] therefore supports the absence of normal diffusion in the easy-axis regime [35, 36]. Other important directions would be to obtain the exact FCS of the SCCA via microscopic computations and to establish a scattering description of thermodynamics, as recently obtained for classical integrable spin chains [37].
@@
Acknowledgements.—We thank Bruno Bertini, Benjamin Doyon, Sarang Gopalakrishnan, Enej Ilievski, Katja Klobas, Marcin Mierzejewski, Peter Prelovšek, Tomaž Prosen, and Herbert Spohn for useful discussions. We are also grateful to Alvise Bastianello for his comments on the manuscript. ŽK is supported by the Simons Foundation as a Junior Fellow of the Simons Society of Fellows (1141511). This paper is dedicated to the memory of Marko Medenjak (1990-2022).

### v3

@@ V Typical fluctuations of deterministic dynamics from Euler hydrodynamics
The assertion that such fluctuations are always transported by the Euler equations regardless of the dynamical exponent stems from the observation that the only source of typical current fluctuations in the SCCA in the single-file limit are the interactions of ballistically-propagating initial typical fluctuations. A similar observation was recently also made for Dirac fluids in Ref. [18]. This is in fact in the same spirit as diffusion from convection, which is known to be the only contribution to the diffusion constants in integrable systems [44, 45, 46] and is generically also present even in non-integrable systems, provided that the eigenvalues of the flux Jacobian $\mathsf{A}$ are non-degenerate [47].
We illustrate the idea at half-filling $b=0$. Since in this case the dynamical exponent of charge transport is $z=2$, the fluctuating fields $\varrho_{i}(x,t)$ need to be rescaled to the diffusive coordinates $\varrho_{i}(x,t)=\hat{\varrho}_{i}(\sqrt{\tau}x,\tau t)$ with the associated currents $j_{i}(x,t)=\sqrt{\tau}\hat{\jmath}_{i}(\sqrt{\tau}x,\tau t)$. The fields $\varrho_{i}(x,t)$ then fluctuate initially as $\varrho_{i}(x,0)\simeq\rho_{i}+\tau^{-1/4}\delta\varrho_{i}(x,0)$, where $\delta\varrho_{i}(x,0)$ is Gaussian distributed as $\langle\delta\varrho_{i}(x,0)\delta\varrho_{j}(y,0)\rangle=\mathsf{C}_{ij}\delta(x-y)$ with the susceptibility matrix $\mathsf{C}_{ij}=\partial\rho_{i}/\partial\beta^{j}$ and $\rho_{i}$ is the background density. The fluctuating particle fields $\delta\varrho_{\pm}(x,0)$ are then transported by the Euler equation $\partial_{t}\delta\varrho_{\pm}(x,t)\pm\sqrt{\tau}\partial_{x}\delta\varrho_{\pm}(x,t)=0$, which can be solved as $\delta\varrho_{\pm}(x,t)\simeq\rho_{\pm}+\tau^{-1/4}\delta\varrho_{\pm}(x\mp\sqrt{\tau}t,0)$. Likewise the charge field $\delta\varrho_{c}(x,t)$ also fluctuates as $\delta\varrho_{c}(x,t)\simeq\tau^{-1/4}\delta\varrho_{c}(x-X_{d}(t),0)$ where $X_{d}(t)$ is the characteristics that satisfies ${\rm d}X_{d}(t)/{\rm d}t=\sqrt{\tau}v(X_{d}(t),t)$. With these we obtain the leading behavior of the fluctuating charge current at the origin $j_{c}(0,t)\simeq\tau^{-1/4}\partial_{t}\int_{0}^{X_{d}(t)}{\rm d}x\,\delta\varrho_{c}(-x,0)$, which in turn allows us to write down the functional integral that describes the contribution to the typical fluctuations by these fields
@@
where $\mathsf{C}^{ij}=[\mathsf{C}^{-1}]_{ij}$ and $Z$ is the normalization factor of the initial probability distribution. The integral is performed exactly in Methods, yielding the distribution
$\mathcal{P}_{\mathrm{typ}}(j)=\frac{1}{2\pi\Delta}\int_{\mathbb{R}}\frac{{\rm d}u}{\sqrt{|u|}}\,e^{-\frac{u^{2}}{2\Delta^{2}}-\frac{j^{2}}{2|u|}}$ (13)
which matches the result reported in Ref. [10]. The breakdown of Gaussianity can be understood by noting which fluctuating fields are responsible for charge current fluctuations. At half-filling, it follows from Eq. (12) that charge fluctuations $\delta\varrho_{c}(x,t)$ that travel along the characteristics $X_{d}(t)$ and cross $x=0$ between $t=0$ and $t=\tau$ contribute to current fluctuations. Since charge fluctuations are transported non-linearly, Gaussianity of inital fluctuations is no longer preserved and typical fluctuation become non-Gaussian. The situation changes dramatically away from half-filling, i.e. for $b\neq 0$ where only particle fluctuations $\delta\varrho_{\pm}(x,t)$ emanating along light cones contribute to charge current fluctuations. Since particle fluctuations are transported linearly, Gaussianity of initial fluctuations is preserved by the dynamics, giving rise to Gaussian typical fluctuations (see Methods for details).
VI Typical fluctuations of stochastic dynamics
The distribution features a cusp at the origin, see right panel of Figure 2. Numerical simulations for $\Gamma>0$ at half-filling indicate that the cusp is rounded off with stochastic crossings while the distribution remains non-Gaussian. This motivates a generalization of the form
@@ VI Typical fluctuations of stochastic dynamics
which reduces to the single-file distribution as $\omega\to 0$ and to a Gaussian distribution (with variance $\omega$) as $\omega\to\infty$. Treating $\omega$ and $\sigma$ as $\Gamma$-dependent fitting parameters, the distribution (14) captures the asymptotic typical distribution for $\Gamma\geq 0$ across 5 order of magnitude as shown in Fig.  2.
VII Conclusions
@@
We have studied anomalous charge current fluctuations in a family of stochastic cellular automata and characterized both large and typical fluctuations based only on Euler hydrodynamics, reproducing the microscopic results obtained in Ref. [10] when specializing to the deterministic limit $\Gamma=0$. The key insight is that initial fluctuations, carried by Euler hydrodynamics, fully control fluctuations both on the typical and large scales in the deterministic case. To implement the idea, we employed BMFT to compute the SCGF and proposed a BMFT-based approach to describe typical fluctuations induced by initial Gaussian fluctuations (in Ref. [48], it is also shown how BMFT provides the full diffusive-order hydrodynamic equation, where the long-range correlations predicted by BMFT are the source of diffusive-order terms). Following Ref. [18], our work further clarifies the hydrodynamic origin of anomalous fluctuations, which we believe is generic among systems that exhibit anomalous fluctuations.
@@
While initial fluctuations remain the only source of large fluctuations even in the stochastic case $\Gamma\neq 0$, it turns out that typical fluctuations are nontrivially influenced by the presence of stochasticity, which is not prescribed by initial fluctuations. We expect that stochasticity induces normal diffusion, and it would be very interesting to reproduce Eq. (14) by extending the formalism we introduced to incorporate normal diffusion. The observation of the single-file distribution in a classical spin chain [8] therefore supports the absence of normal diffusion in the easy-axis regime [49, 50]. It would also be illuminating to obtain the exact FCS of the SCCA with generic $\Gamma$ via microscopic computations and to establish a scattering description of thermodynamics, as recently obtained for classical integrable spin chains [51].
Another important direction to pursue is to apply the BMFT-based framework we have formulated to anomalous typical fluctuations in the isotropic Heisenberg chain. We expect that such an approach would also allow us to describe the hydrodynamic origin of the superdiffusive spin current fluctuations in the model, which have recently been scrutinized extensively.
VIII Acknowledgements
@@
We thank Bruno Bertini, Benjamin Doyon, Sarang Gopalakrishnan, Enej Ilievski, Katja Klobas, Marcin Mierzejewski, Peter Prelovšek, Tomaž Prosen, and Herbert Spohn for useful discussions. We are also grateful to Alvise Bastianello for his comments on the manuscript. ŽK is supported by the Simons Foundation as a Junior Fellow of the Simons Society of Fellows (1141511). This paper is dedicated to the memory of Marko Medenjak (1990-2022).
Appendix A Solution of the BMFT equations
As outlined in the main text, BMFT allows us to write the generating function $\langle e^{\lambda\hat{J}(\tau)}\rangle$ as a functional integral
$\langle e^{\lambda\hat{J}(\tau)}\rangle$ $=\int_{(\mathbb{S})}\mathcal{D}\underline{\varrho}(\cdot,\cdot)e^{-\tau\mathcal{F}[\underline{\varrho}(\cdot,0)]}\delta(\partial_{t}\underline{\varrho}+\partial_{x}\underline{j})$
$=\int_{(\mathbb{S})}\mathcal{D}\underline{\varrho}(\cdot,\cdot)\mathcal{D}\underline{H}(\cdot,\cdot)e^{-\tau S[\underline{\varrho},\underline{H}]},$ (15)
where the rate function $\mathcal{F}[\underline{\varrho}(\cdot,0)]$ reads [37]
with Lagrange multipliers $\beta^{i}(x,0)$, the free energy density $f(x,0)$ associated with the initial fluctuating density profile $\underline{\varrho}(\cdot,0)$, and the background free energy $f$. The long-time behavior of the generating function is thus ruled entirely by the saddle point of the action $S[\underline{\varrho},\underline{H}]$ given by Eq. (8). Such a saddle point is characterized by the saddle point equation $\delta S[\underline{\varrho},\underline{H}]=0$, which can be recast into the following BMFT equations
$\partial_{t}\beta^{\pm}(x,t)\pm\partial_{x}\beta^{\pm}(x,t)\pm b_{\pm}(x,t)\partial_{x}\beta^{c}(x,t)=0,$ (17)
$\partial_{t}H^{\pm}(x,t)\pm\partial_{x}H^{\pm}(x,t)\pm b_{\pm}(x,t)\partial_{x}\beta^{c}(x,t)=0,$ (18)
$\partial_{t}\beta^{c}(x,t)+v(x,t)\partial_{x}\beta^{c}(x,t)=0,$ (19)
$\partial_{t}H^{c}(x,t)+v(x,t)\partial_{x}H^{c}(x,t)=0,$ (20)
with the boundary conditions at $t=0$ and $t=1$
$H^{\pm}(x,0)=\beta^{\pm}(x,0)-\beta^{\pm},$ (21)
$H^{\pm}(x,1)=0,$ (22)
$H^{c}(x,0)=\beta^{c}(x,0)-\beta^{c}+\lambda\Theta(x),$ (23)
$H^{c}(x,1)=\lambda\Theta(x).$ (24)
where $\Theta(x)$ is the Heaviside step function and $b_{\pm}(x,t):=b(x,t)\varrho_{\mp}(x,t)/\varrho(x,t)$.
To solve the BMFT equations, we first start by noting that Eq. (20) can be solved using the method of characteristics as $H^{c}(x,t)=\lambda\Theta(x-x_{2}(t))$, where the characteristic line $x_{2}(t)$ is defined as the solution to ${\rm d}x_{2}(t)/{\rm d}t=v(x_{2}(t),t)$ with $x_{2}(1)=0$. With the boundary condition this then gives another boundary condition $\beta^{c}(x,0)=\beta^{c}+\lambda\left[\Theta(x-x_{2}(t))-\Theta(x)\right]$, which can be used to obtain the full profile of $\beta^{c}(x,t)$
$\beta^{c}(x,t)=\beta^{c}+\lambda\left[\Theta(x-x_{2}(t))-\Theta(x-x_{1}(t))\right],$ (25)
where another characteristic line $x_{1}(t)$ satisfies ${\rm d}x_{1}(t)/{\rm d}t=v(x_{1}(t),t)$ with $x_{1}(0)=0$.
Next we turn to the equations for right/left movers, whose characteristics $x_{\pm}(t)$ are trivial
$x_{\pm}(t)=x\pm t$ (26)
along which the fields $H^{\pm}(x,t)$ evolve according to
$\tfrac{{\rm d}}{{\rm d}t}H^{\pm}(x_{\pm}(x,t)),t)=\mp\lambda b_{\pm}(x_{\pm}(t),t)\delta(x_{\pm}(t)-x_{2}(t)).$ (27)
Integrating the boundary condition $H^{\pm}(x,1)=0$ backwards using Eq. (27), we then obtain $H^{\pm}(x,t)=\lambda b_{2}\left[\Theta(x_{2}(t)-x)-\Theta(\mp(\tau-t)-x)\right]$, where we noted that $b(x_{i}(t),t)=b_{i}$ is constant along $x_{i}(t)$. We emphasize that $b_{2}$ cannot be identified with $\tanh\beta^{c}(0,\tau)$, the value of $\beta^{c}$ at that point being singular due to the incidence of the contour $x_{2}$. However, the solution (18) shows that $H^{\pm}(x,0)$ are constant on the intervals $(-1,x_{2}(0))$ and $(x_{2}(0),1)$ respectively. To determine its value, we note that the pair of Eqs. (18) can be rewritten as $\tfrac{{\rm d}}{{\rm d}t}\left[H^{\pm}(x_{\pm}(t),t)+\log\cosh\beta^{c}(x_{\pm}(t),t)\right]=\mp\lambda b_{1}\delta(t-t_{1})$, which we use to propagate the initial condition (22) from $t=1$ back to $t=0$. With Eq. (18) and the boundary condition Eq. (21) we can now fully specify $\beta^{\pm}(x,0)$. Invoking a similar relation $\tfrac{{\rm d}}{{\rm d}t}\left[\beta^{\pm}(x_{\pm}(t),t)+\log\cosh\beta^{c}(x_{\pm}(t),t)\right]=0$, we are finally in the position to obtain the full profile of $\beta^{\pm}(x,t)$:
$\beta^{\pm}(x,t)=\beta^{\pm}+s[\Theta(\pm t-x)-\Theta(\pm(t-1)-x)$
$\quad-\Theta(x_{1}(t)-x)+\Theta(x_{2}(t)-x)]\frac{\cosh\beta^{c}(\lambda)}{\cosh\beta^{c}},$ (28)
where $\beta^{c}(\lambda):=\beta^{c}+\lambda s$ and $s:={\rm sgn}(x_{1}-x_{2})$. It is a simple matter to argue that the sign $s$ remains constant in time, and is given by $s=\mathrm{sgn}(\lambda b)$. For $\beta^{\pm}=\beta$ and $t\in(0,1)$ the solution is time-independent along the ray $x=0$ and reads
$\beta^{(\lambda),\pm}(0,t)$ $=\beta-2\Theta(\mp\lambda b)\log\left[\frac{\cosh\beta^{c}(\lambda)}{\cosh\beta^{c}}\right].$ (30)
In terms of the solution $\underline{\beta}^{(\lambda)}(x,t)$ of the BMFT equations, the SCGF is computed as [37]
where $\underline{\varrho}^{(\lambda)}(x,t)=\underline{\varrho}[\underline{\beta}^{(\lambda)}(x,t)]$. Plugging Eqs. (29) and (30) into Eq. (31), we obtain Eq. (9). To our knowledge, the above is the first exact solution of full BMFT equations in an interacting model and therefore serves as an independent check of their validity.
Appendix B Comparison between BMFT and BFT
It was argued in Ref. [37] that BMFT in fact reduces to ballistic fluctuation theory (BFT) [42, 52, 53] when the former is specialized to equilibrium. Let us demonstrate that this is indeed the case in the SCCA as well. BFT is a large deviation theory that describes the current fluctuations of the system that supports ballistic transport. A major difference between BFT and BMFT is that BFT is applicable only to systems in equilibrium, which makes the structure of BFT easier to handle. The main observation in BFT is that the insertion of the operator $e^{\lambda\hat{J}(t)}$ in a GGE amounts to a modification of the GGE parameterized by the Lagrange multipliers $\underline{\beta}$ into that parameterized by the $\lambda$-dependent Lagrange multipliers $\underline{\beta}(\lambda)$. Adopting to the charge transport in the SCCA, such multipliers $\underline{\beta}(\lambda)$ satisfy the so-called flow equation
$\partial_{\lambda}\beta^{i}(\lambda)=[\mathrm{sgn}\mathsf{A}(\lambda)]_{c}^{\leavevmode\nobreak\ i},$ (32)
where $\mathsf{A}(\lambda)$ is the flux Jacobian evaluated with respect to the multipliers $\underline{\beta}(\lambda)$. More explicitly, the equations for multipliers read
$\partial_{\lambda}\beta^{\pm}(\lambda)$ =±$=\pm 2b\Theta(\mp(\beta^{+}(\lambda)-\beta^{-}(\lambda))),$ (33)
$\partial_{\lambda}\beta^{c}(\lambda)$ $=\mathrm{sgn}(\beta^{+}(\lambda)-\beta^{-}(\lambda)).$ (34)
In terms of the solution of the flow equation, the SCGF is given by
$F(\lambda)=\int_{0}^{\lambda}{\rm d}\lambda^{\prime}\,j_{c}(\lambda^{\prime}),$ (35)
where $j_{c}(\lambda)$ is the charge current evaluated again with respect to $\underline{\beta}(\lambda)$. The flow equations Eqs. (33) and (34) can be in fact solved explicitly (for $\beta^{\pm}=\beta)$, and the solution turns out to precisely coincide with the solution of the BMFT equations at $x=0$ Eqs. (29) and (30). We thus confirm that the modified $\lambda$-dependent GGE is indeed realized along the ray $x=0$ as proposed in Ref. [37].
Appendix C Typical probability distribution at $\Gamma=0$
In order to evaluate the functional integral Eq. (12), we first need to recast it in terms of normal modes $\delta\varrho=R^{-1}\delta n$ such that $R\mathsf{C}R^{-1}=I$, where the transformation matrix $R^{-1}$ reads
$R^{-1}=\frac{1}{\sqrt{2}}\begin{pmatrix}\Delta_{+}&0&0\\ 0&\Delta_{-}&0\\ b\Delta_{+}&b\Delta_{-}&\sqrt{(1-b^{2})(\rho_{+}+\rho_{-})}\end{pmatrix},$ (36)
with $\Delta_{\pm}=\sqrt{\rho_{\pm}(1-\rho_{\pm})}$. Using this, we can evaluate the normalization factor
$Z$ $=\int_{(\mathbb{R})}\mathcal{D}\underline{\delta\varrho}(\cdot,0)e^{-\frac{\mathsf{C}^{ij}}{2}\int_{\mathbb{R}}{\rm d}x\delta\varrho_{i}(x,0)\delta\varrho_{j}(x,0)}$
$=(2\pi)^{3/2}\det R^{-1}=\sqrt{2\pi(1-b^{2})\rho}\Delta^{2},$ (37)
which allows us to rewrite Eq. (12) as
$\langle e^{\lambda\hat{J}_{c}(\tau)}\rangle$ $\simeq\frac{1}{(2\pi)^{3/2}}\int_{(\mathbb{R})}\mathcal{D}\delta\underline{n}(\cdot,0)e^{-\frac{1}{2}\int_{\mathbb{R}}{\rm d}x(\delta n_{i}(x,0))^{2}}$
×$\quad\times e^{\lambda\tau^{1/4}\sqrt{\rho}\int{\rm d}x\chi(-|X_{d}(1)|<x<0)\delta n_{c}(x,0)]}$
$=\frac{1}{2\pi}\int_{(\mathbb{R})}\mathcal{D}\delta n_{1}(\cdot,0)\mathcal{D}\delta n_{2}(\cdot,0)$
×$\quad\times e^{-\frac{1}{2}\int{\rm d}x[(\delta n_{1}(x,0))^{2}+(\delta n_{2}(x,0))^{2}]}e^{\frac{1}{2}\lambda^{2}\tau^{1/2}\rho|X_{d}(1)|}$ (38)
Assuming $X_{d}(t)$ is monotonic in time, we then have
$\langle e^{\lambda\hat{J}_{c}(\tau)}\rangle$
$\simeq\int_{(\mathbb{R})}\frac{\mathcal{D}\delta n_{1}(\cdot,0)\mathcal{D}\delta n_{2}(\cdot,0)}{2\pi}e^{-\frac{1}{2}\int_{\mathbb{R}}{\rm d}x[(\delta n_{1}(-x,0))^{2}+(\delta n_{2}(x,0))^{2}]}$
×$\quad\times e^{\frac{\Delta\lambda^{2}\tau^{1/4}}{2\sqrt{2}}|\int_{0}^{\sqrt{\tau}}{\rm d}x\,\delta n_{1}(-x,0)-\delta n_{2}(x,0)|}.$ (39)
Introducing the variables $\delta n(x,0)=(\delta n_{1}(-x,0)+\delta n_{2}(x,0)/\sqrt{2}$ and $\delta\bar{n}(x,0)=(\delta n_{1}(-x,0)-\delta n_{2}(x,0))/\sqrt{2}$, we can carry out the integral explicitly
$\langle e^{\lambda\hat{J}_{c}(\tau)}\rangle$ $\simeq\frac{1}{\sqrt{2\pi}}\int_{(\mathbb{R})}\mathcal{D}\delta\bar{n}(\cdot,0)e^{-\frac{1}{2}\int_{\mathbb{R}}{\rm d}x(\delta\bar{n}(x,0))^{2}}$
×$\quad\times e^{\frac{\Delta\lambda^{2}\tau^{1/4}}{2}|\int_{0}^{\sqrt{\tau}}{\rm d}x\,\delta\bar{n}(x,0)|}$
$=e^{\xi^{4}/2}(1+\mathrm{erf}(\xi^{2}/\sqrt{2})),$ (40)
where $\xi^{2}=\Delta\lambda^{2}\sqrt{\tau}/2$ and we invoked the identity $\int_{\mathbb{R}}{\rm d}y\,e^{-x|y|-(y/2)^{2}}=2\sqrt{\pi}e^{x^{2}}\mathrm{erfc}\,x$. This in turn gives the non-Gaussian typical distribution Eq. (13).
The case of $b\neq 0$ can be dealt with in a similar way. Charge transport now has the dynamical exponent $z=1$, thus the initial typical fluctuations are prescribed by the fluctuating fields $\varrho_{i}(x,0)=\hat{\varrho}_{i}(\tau x,0)$
$\varrho_{i}(x,0)\simeq\rho_{i}+\tau^{-1/2}\delta\varrho_{i}(x,0),$ (41)
where $\delta\varrho_{i}(x,0)$ is normal distributed with the susceptibility $\mathsf{C}_{ij}$ as before. Since such fluctuations are disseminated by the Euler equation, the left and right movers remain diffusively fluctuate
$\varrho_{\pm}(x,t)\simeq\rho+\tau^{-1/2}\delta\varrho_{\pm}(x\mp t,0).$ (42)
We thus realize that the Euler-scale charge current $j_{c}(0,t)$ has the leading behavior
$j_{c}(0,t)$ $=b(0,t)\frac{\varrho_{+}(0,t)-\varrho_{-}(0,t)}{2}$
$\simeq b\tau^{-1/2}\frac{\delta\varrho_{+}(-t,0)-\delta\varrho_{-}(t,0)}{2},$ (43)
where in the second line we used Eq. (42). Note that the fluctuation of the normal mode $b(x,t)$ does not contribute here. With these, we can write down the cumulant generating function as a Gaussian integral
×$\quad\times e^{\frac{b\lambda\sqrt{\tau}}{2}\int_{0}^{1}{\rm d}t\,(\delta\varrho_{+}(-t,0)-\delta\varrho_{-}(t,0))}$
$=\frac{1}{2\pi}\int_{(\mathbb{R})}\mathcal{D}\underline{\delta n}(\cdot,0)\,e^{-\int_{\mathbb{R}}{\rm d}x\frac{1}{2}(\delta n_{i}(x,0))^{2}}$
×$\quad\times e^{-\int_{\mathbb{R}}(\mu_{1}(x)\delta n_{1}(x,0)+\mu_{2}(x)\delta n_{2}(x,0))}$ (44)
where
$\mu_{1}(x)=\mu\chi(-1<x<0),\quad\mu_{2}(x)=-\mu\chi(0<x<1)$ (45)
with $\chi(x)$ the indicator function and $\mu=\lambda b\sqrt{\tau}\Delta/2$. Performing the integral explicitly gives $\langle e^{\lambda\hat{J}(\tau)}\rangle\simeq e^{\mu^{2}/2}$. Transforming back to the probability distribution $\mathcal{P}(j\tau^{-1/2z}|\tau)=(2\pi)^{-1}\int_{\mathbb{R}}{\rm d}\lambda\,e^{-{\rm i}\lambda j\tau^{-1/2z}}\langle e^{{\rm i}\lambda J(\tau)}\rangle$, we finally get
$\mathcal{P}^{[0]}_{\mathrm{typ}}(j)=\frac{1}{\sqrt{2\pi\sigma^{2}}}e^{-\frac{j^{2}}{2\sigma^{2}}},\quad\sigma^{2}=(b\Delta)^{2}.$ (46)
As we discussed in the main text, we can make sense of the markedly different behaviors in the two cases $b\neq 0$ and $b=0$ by appreciating which fluctuating fields are responsible for the charge current fluctuations. First, when $b\neq 0$, it can be readily seen in Eq. (44) that the charge current fluctuations are completely determined by the initial particle fluctuations $\delta\varrho_{\pm}$. Since particle fluctuations are propagated freely in time (see Fig. 3a), the initial Gaussian fluctuations remain intact and as a result the typical fluctuations keep Gaussianity. This is in stark contrast to what happens at half-filling where the current fluctuations are now given solely by the charge fluctuations that are transported nonlinearly with the velocity $v$ (see Fig. 3b) according to Eq. (13). Gaussianity is therefore not preserved during the dynamics, resulting in the non-Gaussian distribution.
Figure 3: Initial fluctuations that influence the charge current fluctuations. (a): when $b\neq 0$, the particle fluctuations $\delta\varrho_{\pm}(x,0)$ that cross $x=0$ between $t=0$ and $t=\tau$ give rise to the current fluctuations. (b): when $b=0$ the charge fluctuations $\delta\varrho_{c}(x,0)$ that cross $x=0$ between $t=0$ and $t=\tau$ contribute to the current fluctuations. Note that while the trajectory is now nonlinear, the slope cannot exceed $1$ because $|{\rm d}X_{d}(t)/{\rm d}t|<1$. Here we assumed $X_{d}(t)>0$ for $t\in[0,\tau]$.